solve: 28x – 6(4x – 0.325) = 3x + 0.085

Answer by chitra(359)

The given expression is:


28x – 6(4x – 0.325) = 3x + 0.085


28x - 24x + 1.95 = 3x + 0.085


4x + 1.95 = 3x + 0.085


4x - 3x = 0.085 - 1.95

x = -1.865


Hence, the solution.